OpenStudy (anonymous):

Leslie35: find the simple interest and the total amount due on the loan. Principle $17,400 rate is 3 1/2% Time 6 months what is the simple interest? PLEASE HELP PLEASE

OpenStudy (anonymous):

SI = PRT/100 = 17400 * 3.5 * 1/200 = 87 * 3.5 = 304.5 dollars Amount= 17400 + 304.5 = 17704.5 dollars

OpenStudy (campbell_st):

use I = Prn I = interest, P = principal ( amount invested) r = interest rate as a decimal n = number of time periods. In your question P = 17400, r = 3.5% per ammun or 1.75% per 6 months as a decimal it is 0.0175. n = 1 91 6 month period) I = 17400 x 0.0175 x 1 = 304.50 the interest earnt is $304.50 Amount due = principal + interest